OpenStudy (anonymous):

What is the pH of a 0.050 M concentration of the strong acid HI? PLEASE HELP ME FIND THE ANSWER I AM LOOKING TO LEARN HOW TO DO THIS

OpenStudy (anonymous):

@Nurali @Frostbite @blues @thomaster @radar Can one of you please show me how to do this?

OpenStudy (frostbite):

For a strong acid we assume it fully dissociates and use the definition of pH: \[\LARGE pH=-\log \left( \left[ H ^{+} \right] \right)\]

OpenStudy (frostbite):

In order words we assume the following reaction happens to the end: \[\Large \sf IH + H _{2}O \to H _{3}O ^{+}+I ^{-}\]

OpenStudy (frostbite):

Just put into the calculator: \[\Large pH=-\log( 0.050)=\]

OpenStudy (frostbite):

Answer below if unsure: http://bit.ly/19CIuEh

OpenStudy (anonymous):

Thank you so much!

OpenStudy (frostbite):

No problem at all.